# Reconstructing Movement from Sparse Samples: Enhanced Spatio-Temporal Matching Strategies for Low-Frequency Data

### 10 Mar 2026 v1

@@ 4 Conclusions and Discussion
This paper aimed at enhancing a Map Matching algorithm known as the ST-Matching both to improve its computational efficiency and the precision of the Map Matching results. The study explored several modifications by taking into account the available features of the data, such as GPS uncertainty and the historical usage patterns of the road network. To achieve this goal, four major modifications were suggested: a dynamic buffer for candidate point search based on GPS uncertainty; a dynamic observation probability function where its parameters change with respect to GPS uncertainty; an entirely redesigned temporal score consists of three penalty functions; and a behavioral function based on historical data. They led to the implementation of two new enhanced version of the ST-Matching algorithm: the modified ST-Matching, and the STB-Matching.
Given the absence of ground truth data, the proposed Map Matching enhancements were evaluated through a novel assessment framework. This original methodology employs a multidimensional set of internal metrics, specifically categorized into efficiency, quality, topology, and speed, thereby providing a robust and comprehensive validation of the algorithm’s performance across diverse operational constraints.

### 11 Mar 2026 v2

@@ 4 Conclusions and Discussion
This paper aimed at enhancing a Map Matching algorithm for low frequency data, known as the ST-Matching, both to improve its computational efficiency and the precision of the results. The study explored several modifications by taking into account the available features of the data, such as GPS uncertainty and the historical usage patterns of the road network. To achieve this goal, four major modifications were suggested: a dynamic buffer for candidate point search based on GPS uncertainty; a dynamic observation probability function where its parameters change with respect to GPS uncertainty; an entirely redesigned temporal score consists of three penalty functions; and a behavioral function based on historical data. They led to the implementation of two new enhanced version of the ST-Matching algorithm: the modified ST-Matching, and the STB-Matching.
Given the absence of ground truth data, the proposed Map Matching enhancements were evaluated through a novel assessment framework. This original methodology employs a multidimensional set of internal metrics, specifically categorized into efficiency, quality, topology, and speed metrics, thereby providing a robust and comprehensive validation of the algorithm’s performance across diverse operational constraints.